In Andalusian fowl, a single gene determines feather color. Individuals are either white (aa), gray (Aa), or black (AA). The table gives the observed numbers for each phenotype in a wild population of 2,000 birds. Calculate actual allele frequencies, and the three expected phenotypes if the population is in Hardy Weinberg equilibrium for this locus. Show your work in a way that Phenotype Black (AA) Gray (Aa) White (aa) I can follow your calculations. Draw boxes around each Number 940 of the three final answers. 1055
Answer #1

Answer:

940 AA = 1880 A

1055 Aa = 1055 A & 1055a

5aa = 10 a

Total alleles = 4000

Total alleles, A = 1880 + 1055 = 2935

Frequency of allele, A = 2935 / 4000 = 0.73

Total alleles, a = 10 + 1055 = 10655

Frequency of allele, a = 1065 / 4000 = 0.27

The expected genotype, AA frequency = 0.73 * 0.73 = 0.5329

Number of AA individuals = 0.5329 * 2000 = 1066

The expected genotype, Aa frequency = 2* 0.73 * 0.27 = 0.3942

Number of Aa individuals = 0.3942 * 2000 = 788

The expected genotype, aa frequency = 0.27 * 0.27 = 0.0729

Number of Aa individuals = 0.0729 * 2000 = 146